In excersise 33-41,Use the given information to find the given amount.
38. Wholesale price: $88
Markup Percent: 50%
Any help with this equation would be Greatly appreciated.

Answer by checkley71(8403)

88+.5*88=88+44=$132 ANSWER FOR THE FINAL PRICE.